# EgoCross: Benchmarking Multimodal Large Language Models for Cross-Domain Egocentric Video Question Answering

### v1

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have significantly pushed the frontier of egocentric video question answering (EgocentricQA). However, existing benchmarks and studies are mainly limited to common daily activities such as cooking and cleaning. In contrast, real-world deployment inevitably encounters domain shifts, where target domains differ substantially in both visual style and semantic content. To bridge this gap, we introduce EgoCross, a comprehensive benchmark designed to evaluate the cross-domain generalization of MLLMs in EgocentricQA. EgoCross covers four diverse and challenging domains, including surgery, industry, extreme sports, and animal perspective, representing realistic and high-impact application scenarios. It comprises approximately 1,000 QA pairs across 798 video clips, spanning four key QA tasks: prediction, recognition, localization, and counting. Each QA pair provides both OpenQA and CloseQA formats to support fine-grained evaluation. Extensive experiments show that most existing MLLMs, whether general-purpose or egocentric-specialized, struggle to generalize to domains beyond daily life, highlighting the limitations of current models. Furthermore, we conduct several pilot studies, e.g., fine-tuning and reinforcement learning, to explore potential improvements. We hope EgoCross and our accompanying analysis will serve as a foundation for advancing domain-adaptive, robust egocentric video understanding. Data and codes will be released at: https://github.com/MyUniverse0726/EgoCross.
@@ 1 Introduction
Egocentric videos, which capture how humans perceive and interact with the physical world from a first-person perspective, offer a rich and unique source of data for modeling human behaviors. Understanding egocentric vision is therefore highly valuable for applications such as embodied AI, wearable assistants, and human-to-robot learning. Among various egocentric tasks, video question answering [1, 2], which reflects a model’s understanding through human-like question-answering pairs, has emerged as a particularly challenging yet impactful problem.
Relative early efforts such as EgoVQA [3], EgoTaskQA [4], and EgoSchema [5] have laid the groundwork for egocentric video question answering (EgocentricQA) by introducing some of the first dedicated benchmarks. The rapid progress of Multimodal Large Language Models (MLLMs) has further significantly advanced this field in both benchmark construction and model development. On the benchmark side, EgoThink [6], EgoTempo [7], and EgoTextVQA [8] have been proposed, targeting different aspects of the QA task. On the modeling side, a number of MLLMs specifically designed or adapted for egocentric video understanding have also emerged. Notable examples include EgoVLPv2 [9] and EgoGPT [10], which extend general-purpose MLLMs for EgocentricQA by training on specialized egocentric data.
@@
To answer this question, we introduce EgoCross, a comprehensive benchmark designed to evaluate the cross-domain generalization capabilities of MLLMs in EgocentricQA. EgoCross is built upon three core design principles: (1) emphasis on cross-domain properties, (2) relevance to practical applications, and (3) fine-grained, multi-dimensional model assessment. Following these principles, we carefully curated video sources and developed corresponding QA pairs to reflect real-world, high-impact use cases. Specifically, we selected surgery, industry, extreme sports, and animal perspective, as the four basic domains of our benchmark. These domains exhibit substantial visual and semantic deviations from typical daily-life scenarios, thus posing unique challenges for model generalization. Based on these video sources, we designed a structured data curation pipeline to construct QA pairs across four fundamental QA task types: identification, localization, prediction, and counting, further spanning a total of 15 specific subtasks. To support both discriminative and generative evaluation protocols, each QA instance is annotated in both CloseQA (multiple-choice) and OpenQA (free-form answer) formats. In total, EgoCross consists of approximately 1,000 QA pairs across 798 egocentric video clips, forming a carefully constructed dataset that enables systematic evaluation of cross-domain generalization in EgocentricQA. A visual overview and representative examples are provided in Figure 1.
@@ 2.1 Egocentric Video Understanding
Egocentric video understanding has gained growing attention for modeling human perception and interaction from a first-person view. Beyond perception tasks like action recognition [11], pose estimation [12], and ego-exo correspondence [13, 14], EgocentricQA [3, 4, 5, 6, 7, 8, 15] has emerged as a particularly informative yet challenging task. Widely used EgocentricQA benchmarks include EgoVQA [3], EgoTaskQA [4], and EgoSchema [5], with several new datasets recently introduced to address different challenges. For example, EgoThink [6] focuses on complex reasoning, EgoTempo [7] on temporal understanding, and EgoTextVQA [8] on understanding text from egocentric scenes. The increasing data has also spurred the development of specialized models for egocentric video understanding, typically adapted from MLLMs. However, most existing work remains confined to daily-life scenarios, with limited attention paid to domain shifts. In contrast, our work fills this gap by emphasizing the importance of real-world, out-of-distribution targets and introducing the first comprehensive cross-domain testbed for EgocentricQA.
@@ 2.2 MLLMs for Video Understanding
Recent advancements in Multimodal Large Language Models (MLLMs) have showcased remarkable capabilities in video understanding. General MLLMs such as GPT-4.1 [16], Gemini 2.5 Pro [17], Qwen2.5-VL [18], and InternVL [19] achieve strong performance across a range of video tasks through extensive multimodal pretraining. In parallel, specialized models like Video-LLaMA3 [20] further improve temporal reasoning via dedicated architectural designs. Several MLLMs have also been tailored specifically for egocentric videos, including EgoVLPv2 [9] and EgoGPT [10]. While these models perform well on third-person videos and egocentric videos from common daily scenarios, their ability to generalize to unfamiliar, domain-specific scenarios remains largely unexamined. In this work, we systematically assess how the current state-of-the-art MLLMs generalize to cross-domain egocentric targets, revealing their limitations and offering in-depth analysis to facilitate future research in this direction.
@@ 2.3 Cross-Domain Generalization
Cross-domain generalization is a broad and long-standing challenge in computer vision. Prior work has investigated it across various tasks, including image classification [21, 22, 23, 24, 25, 26, 27, 28], action recognition [29, 30, 31], and object detection [32, 33, 34, 35, 36], achieving remarkable progress via leveraging domain transfer, data augmentation, and efficient fine-tuning techniques. However, these efforts have primarily focused on third-person viewpoints and low-level perception tasks. In egocentric video understanding, domain shifts are particularly pronounced due to drastic variations in scenes, task semantics, and camera motion. A closely related work [37] explores cross-domain few-shot recognition in egocentric videos, but it remains limited to low-level perception tasks. In contrast, EgoCross is the first benchmark specifically designed to evaluate cross-domain generalization in EgocentricQA, tackling both domain gap and high-level reasoning.
@@ 3 EgoCross Dataset
In this section, we provide a comprehensive introduction to the EgoCross benchmark. We begin by discussing the selection of domains, video sources, and the taxonomy of question-answering tasks, followed by an explanation of the data curation pipeline, and conclude with dataset statistics.
@@ 3.1 Source Selection and Task Taxonomy
We established key principles for domain and dataset selection, as well as question-answering task taxonomy: (1) Emphasis on Cross-Domain Properties. We need to select domains with distinct knowledge structures, terminologies, and interactions that differ significantly from everyday scenarios, ensuring the models are challenged by unfamiliar concepts. (2) Impact on Practical Applications. Datasets closely related to real-world applications, e.g., healthcare and industrial operations, are encouraged, as they are expected to foster progress toward practical applications of EgocentricQA. (3) Fine-grained Multi-dimensional Model Assessment. Tasks should span a broad range, covering diverse examination types, such as complex reasoning and spatiotemporal dependencies, and also with comprehensive evaluation metrics.
@@
Based on the above criteria, we select four professional domains that present distinct challenges and high real-world relevance: surgery, industry, extreme sports, and animal perspective. For each, we curated one or two high-quality, open-source datasets with expert-provided meta annotations, each presenting unique perceptual, cognitive, and reasoning demands. The selected domains and their corresponding datasets are detailed as follows:
  • Surgery. The surgical domain represents a highly structured, knowledge-intensive scenario where precision, sequential understanding, and risk-awareness are paramount. To enrich visual diversity, we include two datasets: EgoSurgery [38], which records the videos of open-heart surgeries from the surgeon’s perspective, with fine-grained annotations of hand-tool interactions and surgical phases; and CholecTrack20 [39], which offers laparoscopic videos of cholecystectomy procedures from a tool-centered perspective.
  • Industry. Complex workflows in industrial scenarios demand not only perception of fine object manipulations but also reasoning over procedural sequences and tool-usage logic. We choose ENIGMA-51 [40], a dataset containing real circuit board repair tasks.
  • Extreme Sports. Extreme sports pose unique challenges, such as rare environments, rapid camera motion, and blur, which could well test models’ spatiotemporal perception and high-speed situational reasoning. We include the ExtremeSportFPV [41], which features first-person videos of various extreme sports, including mountain biking, skiing, and skydiving.
  • Animal Perspective. To challenge anthropocentric bias in existing models, we introduce the animal perspective, introducing new motion patterns, camera angles, and semantic focus to the models. EgoPet [42], a dataset featuring egocentric views from animals such as dogs, cats, eagles, and turtles, is thus included.
@@
Following Principle (3), we aim to construct diverse QA pairs to comprehensively assess model capabilities. As illustrated in Figure 3, our evaluation framework is built around four core task categories: Identification, Localization, Prediction, and Counting. Tailored to address the unique challenges of each domain, we further decompose these four broad categories into 15 specific sub-tasks, collectively forming a comprehensive evaluation framework.
  • Identification. Identification tasks evaluate a model’s ability to recognize objects, actions, and events within a video. These tasks require domain-specific knowledge and adaptation to subtle differences in object properties or actions across contexts.
  • Localization. Localization tasks assess a model’s ability to identify the precise spatial or temporal location of objects, actions, or interactions.
  • Prediction. Prediction tasks are designed to test a model’s ability to forecast future actions or outcomes based on the current content.
  • Counting. Counting tasks are proposed to evaluate a model’s ability to track and count distinct instances or occurrences over time.
@@ 3.2 Data Curation Pipeline
Based on the selected data sources and question categories, we developed a multi-stage curation pipeline (Figure 2) with three key stages: meta annotation refinement, QA template design, and batch generation with quality control.
@@
Although the selected datasets provide original annotations, these are typically tailored for simpler, task-specific objectives such as 2D spatial bounding boxes for tool interactions or temporal segments for action classification. To address this, we performed a comprehensive refinement process that involved unifying annotation formats and conducting manual reviews to ensure the label accuracy.
@@
Following the task taxonomy, we manually designed 8 initial QA templates by creating two for each of the four core task categories. To enhance linguistic diversity and complexity, we employed a large language model (Gemini 2.5 pro) to expand the initial templates by generating domain-specific sub-tasks, using the original templates and refined annotations. All LLM-generated questions were then rigorously verified by human annotators to ensure clarity, logical consistency, and answerability based on the video content.
@@
After obtaining the question templates, we perform batch instantiation to generate final QA pairs. For each sampled template, we first randomly extract a corresponding video clips based on its predefined duration, and then derive the ground-truth answer by executing the associated programmatic reasoning over the cropped clips. For comprehensive evaluation, we adopt both the traditional closed-form multiple-choice format (CloseQA) and a more flexible open-ended format (OpenQA) for the answers.
Figure 3: Overview of the EgoCross task taxonomy and statistics. (Top-left) The overall distribution of the four main task categories. (Top-right) The number of questions across target domains. (Bottom) A selection of representative QA examples for each major capability.
@@ 3.3 Dataset Statistics
Our EgoCross benchmark covers four diverse domains: Surgery, Industry, Extreme Sports (XSports), and Animal Perspective (Animal Per.), sourced from five real-world egocentric video datasets. It comprises 798 video clips and 957 QA pairs, spanning 15 sub-task types grouped into four main categories. Table 1 summarizes key statistics of the five datasets, including the number of clips, QA pairs, and average seconds of video durations (Dur.(s)). Figure 3 further illustrates the composition of EgoCross.

### v2

@@ Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have significantly pushed the frontier of egocentric video question answering (EgocentricQA). However, existing benchmarks and studies are mainly limited to common daily activities such as cooking and cleaning. In contrast, real-world deployment inevitably encounters domain shifts, where target domains differ substantially in both visual style and semantic content. To bridge this gap, we introduce EgoCross, a comprehensive benchmark designed to evaluate the cross-domain generalization of MLLMs in EgocentricQA. EgoCross covers four diverse and challenging domains, including surgery, industry, extreme sports, and animal perspective, representing realistic and high-impact application scenarios. It comprises approximately 1,000 QA pairs across 798 video clips, spanning four key QA tasks: prediction, recognition, localization, and counting. Each QA pair provides both OpenQA and CloseQA formats to support fine-grained evaluation. Extensive experiments show that most existing MLLMs, whether general-purpose or egocentric-specialized, struggle to generalize to domains beyond daily life, highlighting the limitations of current models. Furthermore, we conduct several pilot studies, e.g., fine-tuning and reinforcement learning, to explore potential improvements. We hope EgoCross and our accompanying analysis will serve as a foundation for advancing domain-adaptive, robust egocentric video understanding.
Codehttps://github.com/MyUniverse0726/EgoCross
Challengehttps://egocross-benchmark.github.io/
@@ 1 Introduction
Egocentric videos, which capture how humans perceive and interact with the physical world from a first-person perspective, offer a rich and unique source of data for modeling human behaviors. Understanding egocentric vision is therefore highly valuable for applications such as embodied AI, wearable assistants, and human-to-robot learning. Among various egocentric tasks, video question answering (VQA) (Zhong et al. 2022) has emerged as a particularly challenging yet impactful problem. Early efforts like EgoVQA (Fan 2019), EgoTaskQA (Jia et al. 2022), and EgoSchema (Mangalam et al. 2023) laid the groundwork for EgocentricQA by introducing dedicated benchmarks. The rapid progress of Multimodal Large Language Models (MLLMs) has further significantly advanced this field in both benchmark construction and model development. On the benchmark side, EgoThink (Cheng et al. 2024), EgoTempo (Plizzari et al. 2025), and EgoTextVQA (Zhou et al. 2025) have been proposed, targeting different aspects of the QA task. On the modeling side, a number of MLLMs specifically designed or adapted for egocentric video understanding have also emerged. Notable examples include EgoVLPv2 (Pramanick et al. 2023) and EgoGPT (Yang et al. 2025), which extend general-purpose MLLMs for EgocentricQA by training on specialized egocentric data.
@@
To answer this question, we introduce EgoCross, a comprehensive benchmark designed to evaluate the cross-domain generalization capabilities of MLLMs in EgocentricQA. EgoCross is built upon three core design principles: ① emphasis on cross-domain properties, ② relevance to practical applications, and ③ fine-grained, multi-dimensional model assessment. Following these principles, we carefully curated video sources and developed corresponding QA pairs to reflect real-world, high-impact use cases. Specifically, we selected surgery, industry, extreme sports, and animal perspective, as the four basic domains of our benchmark. These domains exhibit substantial visual and semantic deviations from typical daily-life scenarios, thus posing unique challenges for model generalization. Based on these video sources, we designed a structured data curation pipeline to construct QA pairs across four fundamental QA task types: identification, localization, prediction, and counting, further spanning a total of 15 specific subtasks. To support both discriminative and generative evaluation protocols, each QA instance is annotated in both CloseQA (multiple-choice) and OpenQA (free-form answer) formats. In total, EgoCross consists of approximately 1,000 QA pairs across 798 egocentric video clips, forming a carefully constructed dataset that enables systematic evaluation of cross-domain generalization in EgocentricQA. A visual overview and representative examples are provided in Fig. 1.
@@ 2.1 Egocentric Video Understanding
Egocentric video understanding, modeling human perception from a first-person view, has gained growing attention.Beyond basic perception (Wang et al. 2021, 2023), EgocentricQA has emerged as a key task. Initial benchmarks like EgoVQA (Fan 2019), EgoTaskQA (Jia et al. 2022), and EgoSchema (Mangalam et al. 2023) have been joined by new datasets focusing on complex reasoning (Cheng et al. 2024), temporal understanding (Plizzari et al. 2025), and scene text (Zhou et al. 2025). The increasing data has also spurred the development of specialized models for egocentric video understanding, typically adapted from MLLMs. However, most existing work remains confined to daily-life scenarios, with limited attention paid to domain shifts. Our work fills this gap by introducing the first cross-domain testbed for EgocentricQA, emphasizing real-world, out-of-distribution targets.
@@ 2.2 MLLMs for Video Understanding
Recent MLLMs have shown remarkable capabilities in video understanding. General MLLMs such as GPT-4.1 (Achiam et al. 2023), Gemini 2.5 Pro (Comanici et al. 2025), Qwen2.5-VL (Bai et al. 2025), and InternVL (Zhu et al. 2025) achieve strong performance across a range of video tasks through extensive multimodal pretraining. In parallel, specialized models like Video-LLaMA3 (Zhang et al. 2025) further improve temporal reasoning via dedicated architectural designs. Several MLLMs have also been tailored specifically for egocentric video understanding, including EgoVLPv2 (Pramanick et al. 2023) and EgoGPT (Yang et al. 2025). While these models perform well on third-person videos and egocentric videos from common daily scenarios, their ability to generalize to unfamiliar, domain-specific scenarios remains largely unexamined. In this work, we systematically assess how well the current state-of-the-art MLLMs generalize to cross-domain egocentric targets, revealing their limitations and offering in-depth analysis to facilitate future research in this direction.
@@ 2.3 Cross-Domain Generalization
Cross-domain generalization is a broad and long-standing challenge in computer vision. Prior work has investigated it across various tasks, including image classification (Zhu et al. 2019; Fu et al. 2023; Zhao et al. 2020), object detection (Zheng et al. 2020; Li et al. 2025; Zhang et al. 2022), and action recognition (Pan et al. 2020; Xu et al. 2022; Chen et al. 2021a), often leveraging domain transfer, data augmentation, efficient fine-tuning, and meta-learning techniques (Chen et al. 2021b). However, these efforts have primarily focused on third-person viewpoints and low-level perception tasks. In egocentric video understanding, domain shifts are particularly pronounced due to drastic variations in scenes, task semantics, and camera motion. While some works explore cross-domain few-shot recognition in egocentric videos (Hatano et al. 2024) or leverage large models for few-shot knowledge transfer (Ge et al. 2023), these remain limited to low-level perception tasks or more general settings. In contrast, EgoCross is the first benchmark specifically designed to evaluate cross-domain generalization in EgocentricQA, tackling both domain gap and high-level reasoning challenges.
@@ 3 EgoCross Dataset
In this section, we provide a comprehensive introduction to the EgoCross benchmark. We begin by discussing the selection of domains, video sources, and the taxonomy of question-answering tasks, followed by an explanation of the data curation pipeline, and conclude with it dataset statistics.
@@ 3.1 Source Selection and Task Taxonomy
We established key principles for domain and dataset selection, as well as question-answering task taxonomy: ① Emphasis on Cross-Domain Properties. We need to select domains with distinct knowledge structures, terminologies, and interactions that differ significantly from everyday scenarios, ensuring the models are challenged by unfamiliar concepts. ② Impact on Practical Applications. Datasets closely related to real-world applications, e.g., healthcare and industrial operations, are encouraged, as they are expected to foster progress toward practical applications of EgocentricQA. ③ Fine-grained Multi-dimensional Model Assessment. Tasks should span a broad range, covering diverse examination types, such as complex reasoning and spatiotemporal dependencies, and also with comprehensive evaluation metrics.
@@
Based on the above criteria, we select four professional domains that present distinct challenges and high real-world relevance: surgery, industry, extreme sports, and animal perspective. For each, we curated one or two high-quality, open-source datasets with expert-provided meta annotations, each presenting unique perceptual, cognitive, and reasoning demands. The selected domains and the corresponding datasets are as follows:
  • Surgery. The surgical domain represents a highly structured, knowledge-intensive scenario where precision, sequential understanding, and risk-awareness are paramount. To enrich visual diversity, we include two datasets: EgoSurgery (Fujii et al. 2024), which records the videos of open-heart surgeries from the surgeon’s perspective, with fine-grained annotations of hand-tool interactions and surgical phases; and CholecTrack20 (Nwoye et al. 2025), which offers laparoscopic videos of cholecystectomy procedures from a tool-centered perspective. This dataset offers a rare yet insightful egocentric perspective captured from a tool rather than a typical human operator.
  • Industry. Complex workflows in industrial scenarios demand not only perception of fine object manipulations but also reasoning over procedural sequences and tool-usage logic. We choose ENIGMA-51 (Ragusa et al. 2024), a dataset containing real circuit board repair tasks.
  • Extreme Sports. Extreme sports pose unique challenges, such as rare environments, rapid camera motion, and blur, which could well test models’ spatiotemporal perception and high-speed situational reasoning. We include the ExtremeSportFPV (Singh et al. 2017), which features first-person videos of various extreme sports, including mountain biking, skiing, and skydiving.
  • Animal Perspective. To challenge anthropocentric bias in existing models, we introduce the animal perspective, introducing new motion patterns, camera angles, and semantic focus to the models. EgoPet (Bar et al. 2024), a dataset featuring egocentric views from animals such as dogs, cats, eagles, and turtles, is thus included.
The selected domains and video sources align well with our principles ① and ②.
@@
Following Principle ③, we aim to construct diverse QA pairs to comprehensively assess model capabilities. As illustrated in Fig. 3, our evaluation framework is built around four core task categories: Identification, Localization, Prediction, and Counting. Tailored to address the unique challenges of each domain, we further decompose these four broad categories into 15 specific sub-tasks, collectively forming a comprehensive evaluation framework. In the following, we provide an overview of the 4 core tasks. Detailed sub-tasks and representative examples can be found in Fig. 3 and Appendix.
  • Identification. Identification tasks evaluate a model’s ability to recognize objects, actions, and events within a video. These tasks require domain-specific knowledge and adaptation to subtle differences in object properties or actions across contexts. For example, in surgical scenarios, instruments like forceps and scissors share similar shapes and colors, placing high demands on the models.
  • Localization. Localization tasks assess a model’s ability to identify the precise spatial or temporal location of objects, actions, or interactions. These tasks require the model to understand spatiotemporal relationships and adjust to the variations in object positioning and motion patterns across different environments. In industrial assembly, for example, locating tools or components in a cluttered workspace is particularly challenging due to partial occlusions and rapid movements of small objects.
  • Prediction. Prediction tasks test a model’s ability to forecast future actions or outcomes based on the current content. Thus, models are expected to grasp the underlying procedural or causal relationships in unseen domains. For example, in surgery, predicting the next phase of the procedure, such as transitioning from suturing to wound closure, relies on models’ understanding of established surgical patterns, which can vary significantly across different domains like industrial assembly or extreme sports.
  • Counting. Counting tasks evaluate a model’s ability to track and count distinct instances or occurrences over time. These tasks require precise identification and temporal aggregation of visual elements, which becomes increasingly complex when dealing with fast-paced or dynamic scenarios. In extreme sports, for example, counting the number of tricks or jumps requires tracking high-velocity actions and differentiating between overlapping movements in a dynamic, cluttered environment.
@@ 3.2 Data Curation Pipeline
Based on the selected data sources and question categories, we developed a multi-stage curation pipeline (Fig. 2) with three key stages: meta annotation refinement, QA template design, and batch generation with quality control. The detailed procedures are described below.
@@
Although the selected datasets provide original annotations, these are typically tailored for simpler, task-specific objectives such as 2D spatial bounding boxes for tool interactions or temporal segments for action classification. In addition, the annotation formats vary significantly between datasets. To address this, we performed a comprehensive refinement process that involved unifying annotation formats and conducting manual reviews to ensure the label accuracy. This refinement step was essential for constructing a reliable ground truth, which serves as the foundation for all subsequent QA generation.
@@
Figure 3: An overview of the EgoCross task taxonomy and statistics. (Top-left) The overall distribution of the four main task categories: Identification, Localization, Prediction, and Counting. (Top-right) The number of questions across the four primary domains. (Bottom) A selection of representative QA examples for each major capability is presented. For a more comprehensive list of examples, please see the Appendix.
@@
Following the task taxonomy, we manually designed 8 initial QA templates by creating two for each of the four core task categories. To enhance linguistic diversity and complexity, we employed a large language model (Gemini 2.5 pro) to expand the initial templates by generating domain-specific sub-tasks, using the original templates and refined annotations. All LLM-generated questions were then rigorously verified by human annotators to ensure clarity, logical consistency, and answerability based on the video content. Annotators also specified the programmatic reasoning steps and the expected answer format (e.g., object names, timestamps) for each template to ensure that the questions are both challenging and solvable.
@@
After obtaining the question templates, we perform batch instantiation to generate final QA pairs. For each sampled template, we first randomly extract a corresponding video clips based on its predifined duration, and then derive the ground-truth answer by executing the associated programmatic reasoning over the cropped clips. For comprehensive evaluation, we adopt both the traditional closed-form multiple-choice format (CloseQA) and a more flexible open-ended format (OpenQA) for the answers. In CloseQA, each question is accompanied by one correct answer and three distractors randomly sampled from the same answer type. For OpenQA, the answer consists of the full reasoning steps, which is further refined by a LLM. To ensure the data quality at scale, we conducted a final quality control check by randomly sampling and verifying 10% of QA pairs from each domain.
@@ 3.3 Dataset Statistics
Our EgoCross benchmark covers four diverse domains: Surgery, Industry, Extreme Sports (XSports), and Animal Perspective (Animal Per.), sourced from five real-world egocentric video datasets. It comprises 798 video clips and 957 QA pairs, spanning 15 sub-task types grouped into four main categories. Tab.1 summarizes key statistics of the five datasets, including the number of clips, QA pairs, and average seconds of video durations (Dur.(s)). Fig.3 further illustrates the composition of EgoCross, including: the distribution of the four primary QA task categories, the 15 sub-task types, the question counts across domains, and representative QA examples for each major capability.